# On geometrical origin of Kodama vector

### 26 Feb 2024 v1

@@ I Introduction and summary
Recently, it is shown that in three-dimensional axisymmetric spacetimes even for non-warped product spacetimes such as rotating ones, the notion of the Kodama vector can be extended [5, 6]. This vector can provide a local conserved current and quasilocal mass taking into account of angular momentum, as in the cases of warped product spacetimes. This fact suggests that warped product does not seem to be necessary to admit the Kodama vector.
In this paper we show properties of the Kodama vector geometrically originate from conformal Killing-Yano (CKY) two-form. Various conserved currents and charges associated with (conformal) Killing tensors and (conformal) Killing-Yano forms have been reported in the literature [7, 8, 9, 10, 11, 14, 15, 12, 13]. What we emphasis here is that the Kodama vector is the so-called associated vector with a CKY two-form. In particular, all the well-known spacetimes admitting the Kodama vector have closed conformal Killing-Yano (CCKY) two-form.
Furthermore, we show that the associated vector of the CKY two-form can yield conserved currents not only for the Einstein tensor but also for each order of the Lovelock tensor [16, 17]. This means that the Kodama vector provides a locally conserved energy current in Lovelock gravity, which has been partially proved and conjectured for symmetric spacetimes such as spherically symmetric one in [18, 19].
@@ II.1 CKY two-form and conserved current for Einstein tensor
where the vector field $K_{a}$ is the so-called associated vector of $h_{ab}$. If $L_{abc}=0$, $h_{ab}$ reduces to a closed conformal Killing-Yano (CCKY) two-form. In this case, a Hodge dual of $h_{ab}$ yields a Killing-Yano $(D-2)$-form $f_{a_{1}\cdots a_{D-2}}$ which satisfies $\nabla_{a}f_{b_{1}\cdots b_{D-2}}=\nabla_{[a}f_{b_{1}\cdots b_{D-2}]}$.
@@
Thus, the associated vector $K^{a}$ for a conformal Killing-Yano two-form $h_{ab}$ provides the same properties as Kodama vectors and $G_{ab}K^{b}$ becomes a locally conserved current. [^1] These properties have been pointed out in Refs. [10, 14, 13], where $G_{ab}K^{b}$ is referred to as “Einstein current.” We note that if the spacetime is an Einstein space, i.e., $R_{ab}=\lambda g_{ab}$, then Eq. (4) leads to the Killing equation $\nabla_{a}K_{b}+\nabla_{b}K_{a}=0$ [21]. This implies a version of Birkhoff’s theorem that the Kodama vector becomes the Killing vector in vacuum with a cosmological constant. In four dimensions, the relation between CKY two-form and Birkhoff’s theorem was discussed [23].
@@
Since this current is given by a divergence of two-form “potential” in Eq. (7), we can explicitly see the current is locally conserved. It is worth noting that the expressions in the first and second lines of (7) is valid in $D>3$ dimensions, because both $\mathcal{P}_{abcd}\equiv R_{abcd}-2R_{a[c}g_{d]b}+2R_{b[c}g_{d]a}+Rg_{a[c}g_{d]b}$ [^2] This rank-$4$ tensor is divergence free and its indices have the same symmetries of the Riemann tensor, which can be also written as $\delta_{abb_{1}b_{2}}^{cda_{1}a_{2}}R_{a_{1}a_{2}}{}^{b_{1}b_{2}}=4\mathcal{P}_{ab}{}^{cd}$ by using the generalized Kronecker delta symbol. This type of tensor has been used in Ref. [31], for example. and $W_{abcd}$ are identically zero in three dimensions. However, that in the last line is valid even in $D=3$ dimensions. We note that $C_{abc}h^{bc}$ is a so-called Cotton current in Ref. [11]. [^3] A conserved current for the Cotton tensor was discussed in Ref. [24], also.
In a specific case, if $h_{ab}$ is a Killing-Yano tensor, then the two-form field $R_{abcd}h^{cd}+4R_{[a}{}^{c}h_{b]c}+Rh_{ab}$ can be conserved. This is referred to as the Yano current [8]. It is equivalent to the fact that the associated vector for the Killing-Yano tensor will vanish in Eq. (7).
@@ II.2 Generalization to Lovelock tensor
The second equality follows from the second Bianchi identity, $\nabla_{[a}R_{bc]de}=0$, and the third equality does from the first Bianchi identity. Since $F_{ab}$ is anti-symmetric, $G^{(n)}{}_{ad}K^{d}$ is divergence free. Hence, we have also local conserved currents for the $n$th order Lovelock tensor as
@@
Note that, for $n=1$, the previous result for the Einstein tensor is obviously reproduced. On arbitrary spacelike hypersurfaces $\Sigma$ with a common boundary $\partial\Sigma$, by using Stokes’ theorem, we have a conserved charge written in the boundary integral. A $n$th order quasilocal charge becomes
@@
We note that the potential two-form field (11) seems to be very similar to a part of the Killing-Lovelock potential [26, 27] to define improved Komar integrals in Lovelock theory. The $n$th Killing-Lovelock potential for the $n$th order Lovelock term, however, consists of ($n-1$) powers of the Riemann tensor. On the other hand, in Ref [15], the authors introduced two-form field with the same powers of the Riemann tensor as (11) for Killing-Yano two-form but not for conformal Killing-Yano two-form. In that case, the two-form itself is conserved.
@@ III.1 Warped-product spacetime
It is known that warped-product spacetimes with two-dimensional base possess the Kodama vector field. We revisit the known results for the Kodama vector in terms of the CKY two-form (also see appendix $D$ in [28]).
@@
where ${}^{(\gamma)}\epsilon_{ab}$ is $2$-dimensional volume form associated with the metric $\gamma_{\mu\nu}$. Note that this is equivalent to the Hodge dual $f=*h$ being a Killing-Yano ($D-2$)-form.
@@
where the conventional Kodama vector is given by $K^{a}=-{}^{(\gamma)}\epsilon^{ab}\nabla_{b}r$. In fact, the warp factor $r$ is given by a “norm” of the CCKY two-form $h$ [or KY ($D-2$)-form $f$] as
@@
If we consider the Einstein equation with a cosmological constant term, $G_{ab}+\Lambda g_{ab}=T_{ab}$, a quasilocal mass
@@
is obtained by combining two conserved charges, including only the contribution of matter. It is built from the CCKY two-form and the Ricci scalar on the fiber $\mathcal{F}$.
@@ III.2 Three-dimensional spacetime
In three dimensions one can consider spacetimes are not warped product but axisymmetric such as a rotating spacetime with angular momentum. In this case the Kodama vector can be defined and it provides conserved current and charge [5, 6].
@@
This is the extended Kodama vector introduced in [5, 6]. Note that $\nabla_{a}\psi_{b}=\epsilon_{abc}K^{c}$. We have
@@
If we assume that $\psi^{a}$ is an axial Killing vector and the Einstein equation $G_{ab}+\Lambda g_{ab}=T_{ab}$ satisfies, the following scalar functions
@@ III.3 Generalized Misner-Sharp mass in Lovelock gravity
Since, for each order of the Lovelock tensor, each current and each charge are conserved, linear combinations of these quantities should be conserved. Hence, according to the field equations, they can reproduce the generalized Misner-Sharp quasilocal mass in Lovelock gravity, which has been proposed in Ref [18, 19].
@@
The conserved currents associated with Killing vectors and Kodama vectors have similar structure [26, 27] built from the following quantities: $\mathcal{P}^{(n)}{}_{ab}{}^{cd}\equiv\delta^{cda_{1}\cdots a_{2n}}_{abb_{1}\cdots b_{2n}}R_{a_{1}a_{2}}{}^{b_{1}b_{2}}\cdots R_{a_{2n-1}a_{2n}}{}^{b_{2n-1}b_{2n}}$. Because a Killing vector $\xi^{a}$ is divergence free, we obtain a $2$-form potential $\omega_{ab}$ such that $\xi^{a}=\nabla_{b}\omega^{ab}$. A part of Komar-type potential is given by $\mathcal{P}^{(n)}_{abcd}\omega^{cd}$. On the other hand, a Kodama vector is provided by a CKY $2$-form $h_{ab}$ as $K^{a}=-\nabla_{b}h^{ab}/(D-1)$ and the potential is given by $\mathcal{P}^{(n)}_{abcd}h^{cd}$. It is fascinating to explore the relation between these conserved current.
@@ Acknowledgements.
I would like to thank T. Houri and M. Nozawa for valuable comments. This work was supported in part by JSPS KAKENHI Grant No. JP21H05186.
@@ Appendix A Curvature tensors on warped product spacetimes
$=\frac{1}{r^{2n}}\sum_{l=0}^{n}{}_{n}C_{l}\frac{2^{n}(D-2-2n+2l)!}{(D-2-2n)!}(K^{a}K_{a})^{l}\mathcal{L}^{(n-l)}_{\omega},$
@@
$\mathcal{L}_{\omega}^{(k)}\equiv\frac{1}{2^{k}}\delta^{I_{1}\cdots I_{2k}}_{J_{1}\cdots J_{2k}}{}^{(\omega)}R_{I_{1}I_{2}}{}^{J_{1}J_{2}}\cdots{}^{(\omega)}R_{I_{2k-1}I_{2k}}{}^{J_{2k-1}J_{2k}}.$ (41)

### 12 Mar 2024 v2

@@ I Introduction and summary
Recently, it is shown that in three-dimensional axisymmetric spacetimes even for non-warped product spacetimes such as rotating ones, the notion of the Kodama vector can be extended [5, 6]. This vector can provide a local conserved current and quasilocal mass taking into account of angular momentum, as in the cases of warped product spacetimes. This fact suggests that warped product does not seem to be necessary for a spacetime to admit the Kodama vector.
In this paper we show properties of the Kodama vector geometrically originate from conformal Killing-Yano (CKY) two-form. Various conserved currents and charges associated with (conformal) Killing tensors and (conformal) Killing-Yano forms have been reported in the literature [7, 8, 9, 10, 11, 14, 15, 12, 13]. What we emphasis here is that the Kodama vector is the so-called associated vector with a CKY two-form. In particular, all the well-known spacetimes admitting the Kodama vector have closed conformal Killing-Yano (CCKY) two-forms, which belong to a subclass of CKY two-forms.
Furthermore, we show that the associated vector of the CKY two-form can yield conserved currents not only for the Einstein tensor but also for each order of the Lovelock tensor [16, 17]. This means that the Kodama vector provides a locally conserved energy current in Lovelock gravity, which has been partially proved and conjectured for symmetric spacetimes such as spherically symmetric one in [18, 19]. (In warped product spacetimes of a two-dimensional base and an Einstein space, the Kodama vector and the Misner-Sharp quasilocal mass were studied in Ref. [20].)
@@ II.1 CKY two-form and conserved current for Einstein tensor
where the vector field $K_{a}$ is the so-called associated vector of $h_{ab}$. If $L_{abc}=0$, $h_{ab}$ reduces to a closed conformal Killing-Yano (CCKY) two-form. In this case, a Hodge dual of $h_{ab}$ yields a Killing-Yano $(D-2)$-form $f_{a_{1}\cdots a_{D-2}}$, which satisfies $\nabla_{a}f_{b_{1}\cdots b_{D-2}}=\nabla_{[a}f_{b_{1}\cdots b_{D-2}]}$.
@@
Thus, the associated vector $K^{a}$ for a conformal Killing-Yano two-form $h_{ab}$ provides the same properties as Kodama vectors and $G_{ab}K^{b}$ becomes a locally conserved current. [^1] These properties have been pointed out in Refs. [10, 14, 13], where $G_{ab}K^{b}$ is referred to as “Einstein current.” We note that if the spacetime is an Einstein space, i.e., $R_{ab}=\lambda g_{ab}$, then Eq. (4) leads to the Killing equation $\nabla_{a}K_{b}+\nabla_{b}K_{a}=0$ [22]. This implies a version of Birkhoff’s theorem that the Kodama vector becomes a Killing vector in vacuum with a cosmological constant. In four dimensions, the relation between CKY two-form and Birkhoff’s theorem was discussed [24].
@@
Since $G_{ab}K^{b}$ is given by a divergence of two-form “potential” in Eq. (7), we can explicitly see this current is locally conserved. It is worth noting that the expressions in the first and second lines of (7) are valid in $D>3$ dimensions, because both $\mathcal{P}_{abcd}\equiv R_{abcd}-2R_{a[c}g_{d]b}+2R_{b[c}g_{d]a}+Rg_{a[c}g_{d]b}$ [^2] This rank-$4$ tensor is divergence free and its indices have the same symmetries of the Riemann tensor, which can be also written as $\delta_{abb_{1}b_{2}}^{cda_{1}a_{2}}R_{a_{1}a_{2}}{}^{b_{1}b_{2}}=4\mathcal{P}_{ab}{}^{cd}$ by using the generalized Kronecker delta symbol. This type of tensor has been used in Ref. [32], for example. and $W_{abcd}$ are identically zero in three dimensions. However, that in the last line is valid even in $D=3$ dimensions. We note that $C_{abc}h^{bc}$ is a so-called Cotton current in Ref. [11]. [^3] A conserved current for the Cotton tensor was discussed in Ref. [25], also.
In a specific case, if $h_{ab}$ is a Killing-Yano tensor, then the “potential” two-form field $R_{abcd}h^{cd}+4R_{[a}{}^{c}h_{b]c}+Rh_{ab}$ itself can be conserved. This is referred to as the Yano current [8]. It is equivalent to the fact that the associated vector for the Killing-Yano tensor will vanish in Eq. (7).
@@ II.2 Generalization to Lovelock tensor
The second equality follows from the second Bianchi identity, $\nabla_{[a}R_{bc]de}=0$, and the third equality does from the first Bianchi identity. Since $F_{ab}$ is anti-symmetric, $G^{(n)}{}_{ad}K^{d}$ is divergence free. Hence, we have also a local conserved current for the $n$th order Lovelock tensor as
@@
Note that, for $n=1$, the previous result for the Einstein tensor is obviously reproduced. On arbitrary spacelike hypersurfaces $\Sigma$ with a common boundary $\partial\Sigma$, by using Stokes’ theorem, we have a conserved charge written in the boundary integral. An $n$th order quasilocal charge becomes
@@
We note that the potential two-form field (11) seems to be very similar to a part of the Killing-Lovelock potential [27, 28] to define improved Komar integrals in Lovelock theory. The $n$th Killing-Lovelock potential for the $n$th order Lovelock term, however, consists of ($n-1$) powers of the Riemann tensor. On the other hand, in Ref [15], the authors introduced a two-form field with the same powers of the Riemann tensor as (11) for Killing-Yano two-forms but not for conformal Killing-Yano two-forms. In that case, the two-form field itself is conserved.
@@ III.1 Warped-product spacetime
It is known that warped-product spacetimes with two-dimensional base possess the Kodama vector field. We revisit the known results for the Kodama vector in terms of CKY two-forms (also see appendix D in [29]).
@@
where ${}^{(\gamma)}\epsilon_{ab}$ is the $2$-dimensional volume form associated with the metric $\gamma_{\mu\nu}$. Note that this is equivalent to the Hodge dual $f=*h$ being the Killing-Yano ($D-2$)-form.
@@
where the conventional Kodama vector is given by $K^{a}=-{}^{(\gamma)}\epsilon^{ab}\nabla_{b}r$. In fact, the warp factor $r$ is given by a “norm” of the CCKY two-form $h$ [or the KY ($D-2$)-form $f$] as
@@
If we consider the Einstein equation with a cosmological constant term, $G_{ab}+\Lambda g_{ab}=T_{ab}$, the Misner-Sharp quasilocal mass
@@
is obtained by combining two conserved charges, including only the contribution of matter without a cosmological constant. It is built from the CCKY two-form and the Ricci scalar on the fiber $\mathcal{F}$.
@@ III.2 Three-dimensional spacetime
In three dimensions one can consider that spacetimes are not warped product but axisymmetric, such as a rotating spacetime with angular momentum. In this case the Kodama vector can be defined and it provides conserved current and charge [5, 6].
@@
This is the extended Kodama vector, which has been introduced in [5, 6]. Note that $\nabla_{a}\psi_{b}=\epsilon_{abc}K^{c}$. We have
@@
If we assume that $\psi^{a}$ is an axial Killing vector and the Einstein equation $G_{ab}+\Lambda g_{ab}=T_{ab}$ is satisfied, the following scalar functions
@@ III.3 Generalized Misner-Sharp mass in Lovelock gravity
Since, for each order of the Lovelock tensor, each current and each charge are conserved, linear combinations of these quantities should be conserved. Hence, according to the field equations, they can reproduce the generalized Misner-Sharp quasilocal mass in Lovelock gravity, which has been proposed in Refs. [18, 19].
We note that, when the ($D-2$)-dimensional subspace is described by Einstein spaces as well as maximally symmetric spaces, the Misner-Sharp quasilocal mass was provided in Ref. [20]. In that case, the quasilocal mass contains the Weyl curvature of the ($D-2$)-dimensional Einstein space. (More generally, it comprises the sum of every order of Lovelock terms for the ($D-2$)-dimensional subspace, as shown in appendix A.)
@@
The conserved currents associated with Killing vectors and Kodama vectors have a similar structure [27, 28] built from the following quantities: $\mathcal{P}^{(n)}{}_{ab}{}^{cd}\equiv\delta^{cda_{1}\cdots a_{2n}}_{abb_{1}\cdots b_{2n}}R_{a_{1}a_{2}}{}^{b_{1}b_{2}}\cdots R_{a_{2n-1}a_{2n}}{}^{b_{2n-1}b_{2n}}$, which are crucial to the Euler-Lagrange equations in Lovelock gravity [34, 35, 26]. Because a Killing vector $\xi^{a}$ is divergence free, we obtain a $2$-form potential $\omega_{ab}$ such that $\xi^{a}=\nabla_{b}\omega^{ab}$. A part of Komar-type potential is given by $\mathcal{P}^{(n)}_{abcd}\omega^{cd}$. On the other hand, a Kodama vector is provided by a CKY $2$-form $h_{ab}$ as $K^{a}=-\nabla_{b}h^{ab}/(D-1)$ and the potential is given by $\mathcal{P}^{(n)}_{abcd}h^{cd}$. It is fascinating to explore the relation between these conserved current.
@@ Acknowledgements.
I would like to thank T. Houri, M. Nozawa, and N. Dadhich for valuable comments. This work was supported in part by JSPS KAKENHI Grant No. JP21H05186.
@@ Appendix A Curvature tensors on warped product spacetimes
$=\frac{2^{n}}{r^{2n}}\sum_{l=0}^{n}{}_{n}C_{l}\frac{(D-2-2n+2l)!}{(D-2-2n)!}(K^{a}K_{a})^{l}\mathcal{R}^{(n-l)}_{\omega},$
@@
$\mathcal{R}_{\omega}^{(k)}\equiv\frac{1}{2^{k}}\delta^{I_{1}\cdots I_{2k}}_{J_{1}\cdots J_{2k}}{}^{(\omega)}R_{I_{1}I_{2}}{}^{J_{1}J_{2}}\cdots{}^{(\omega)}R_{I_{2k-1}I_{2k}}{}^{J_{2k-1}J_{2k}}.$ (41)
@@
If $\omega_{IJ}$ on the fiber $\mathcal{F}$ is a ($D-2$)-dimensional Einstein metric, i.e, ${}^{(\omega)}R_{IJ}=k(D-3)\omega_{IJ}$, then we have ${}^{(\omega)}R_{IJ}{}^{KL}={}^{(\omega)}W_{IJ}{}^{KL}+k\delta_{IJ}^{KL}$, where ${}^{(\omega)}W_{IJKL}$ is the Weyl tensor with respect to $\omega_{IJ}$. Equation (40) reduces to
$\delta^{I_{1}\cdots I_{2n}}_{J_{1}\cdots J_{2n}}R_{I_{1}I_{2}}{}^{J_{1}J_{2}}\cdots R_{I_{2n-1}I_{2n}}{}^{J_{2n-1}J_{2n}}=\frac{2^{n}}{r^{2n}}\sum_{l=0}^{n}{}_{n}C_{l}\frac{(D-2-2n+2l)!}{(D-2-2n)!}(k+K^{a}K_{a})^{l}\mathcal{W}^{(n-l)}_{\omega},$ (43)
where $\mathcal{W}_{\omega}^{(k)}$ has been obtained by replacing the Riemann tensor ${}^{(\omega)}R_{IJ}{}^{KL}$ with the Weyl tensor ${}^{(\omega)}W_{IJ}{}^{KL}$ in Eq. (41).
Appendix B Formulae for curvature polynomials
In this appendix, we summarize basic properties of curvature polynomials in $D$-dimensions. (See, for example, Refs. [26, 34, 35])
The $n$th order Lovelock scalar and Lovelock-Ricci tensor, which are respectively analogous to the Ricci scalar and the Ricci tensor for $n=1$, are
$\mathcal{R}^{(n)}$ $\equiv\frac{1}{2^{n}}\delta^{a_{1}a_{2}\cdots a_{2n}}_{b_{1}b_{2}\cdots b_{2n}}R_{a_{1}a_{2}}{}^{b_{1}b_{2}}\cdots R_{a_{2n-1}a_{2n}}{}^{b_{2n-1}b_{2n}},$ (44)
$\mathcal{R}^{(n)a}{}_{b}$ $\equiv\frac{n}{2^{n}}\delta^{a_{1}a_{2}\cdots a_{2n}}_{bb_{2}\cdots b_{2n}}R_{a_{1}a_{2}}{}^{ab_{2}}\cdots R_{a_{2n-1}a_{2n}}{}^{b_{2n-1}b_{2n}}.$ (45)
The $n$th order Lovelock tensor, which is the analog of the Einstein tensor, is given by
$G^{(n)a}{}_{b}$ $\equiv-\frac{1}{2^{n+1}}\delta^{aa_{1}a_{2}\cdots a_{2n}}_{bb_{1}b_{2}\cdots b_{2n}}R_{a_{1}a_{2}}{}^{b_{1}b_{2}}\cdots R_{a_{2n-1}a_{2n}}{}^{b_{2n-1}b_{2n}}$ (46)
$=\mathcal{R}^{(n)a}{}_{b}-\frac{1}{2}\mathcal{R}^{(n)}g^{a}{}_{b},$
where the last equality is easily verified by using the following formula:
$\delta^{aa_{1}a_{2}\cdots a_{2n}}_{bb_{1}b_{2}\cdots b_{2n}}=g^{a}_{b}\delta^{a_{1}a_{2}\cdots a_{2n}}_{b_{1}b_{2}\cdots b_{2n}}-\sum^{2n}_{k=1}g^{a}_{b_{k}}\delta^{a_{1}a_{2}\cdots a_{k}\cdots a_{2n}}_{b_{1}b_{2}\cdots b\cdots b_{2n}}.$ (47)
An $n$th order rank-$4$ tensor that consists of $n$ powers of the Riemann tensor is given by
$\mathcal{P}^{(n)}{}_{ab}{}^{cd}\equiv\delta^{cda_{1}a_{2}\cdots a_{2n}}_{abb_{1}b_{2}\cdots b_{2n}}R_{a_{1}a_{2}}{}^{b_{1}b_{2}}\cdots R_{a_{2n-1}a_{2n}}{}^{b_{2n-1}b_{2n}}.$ (48)
Its indices have the same properties as those of the Riemann tensor:
$\mathcal{P}^{(n)}{}_{abcd}=-\mathcal{P}^{(n)}{}_{bacd}=-\mathcal{P}^{(n)}{}_{abdc},\quad\mathcal{P}^{(n)}{}_{abcd}=\mathcal{P}^{(n)}{}_{cdab},\quad\mathcal{P}^{(n)}{}_{[abc]d}=0,$ (49)
and, in addition, it is divergence free for each index:
$\nabla^{a}\mathcal{P}^{(n)}{}_{abcd}=0.$ (50)
This tensor has various useful properties as follows. The contraction yields
$\mathcal{P}^{(n)}{}_{ac}{}^{bc}=-2^{n+1}(D-2n-1)G^{(n)}{}_{a}{}^{b}.$ (51)
Furthermore, we have
$\mathcal{R}^{(n)}=\frac{1}{2^{n}}\mathcal{P}^{(n-1)}{}_{abcd}R^{abcd},\quad\mathcal{R}^{(n)a}{}_{b}=\frac{n}{2^{n}}\mathcal{P}^{(n-1)}{}_{bcde}R^{acde}.$ (52)